The width of a rectangle is 3ft. less than the length. The perimeter is 10ft. Find the width
Answer by Earlsdon(6294)

Let L = length and W = width. From the desciption we can write:
W = L-3 or L = W+3
The perimeter of a rectangle is given by:
P = 2L+2W = 10 Substiute the L from above and solve:
2(W+3)+2W = 10 Simplify and solve for W.
2W+6+2W = 10
4W+6 = 10 Subtract 6 from both sides.
4W = 4 Divide both sides by 4.
W = 1
The width is 1 foot.
Check:
P = 2L+2W = 2(W+3) + 2W = 2(1+3) + 2(1) = 2(4) + 2 = 10 Good!